OpenStudy (anonymous):

Write as a simplified single angle expressions of either sine or cosine. cos(pi/7)cos(pi/5)-sin(pi/7)sin(pi/5)

OpenStudy (anonymous):

this is the addition angle formula for cosine

OpenStudy (anonymous):

so your job is to add up \[\frac{\pi}{7}+\frac{\pi}{5}\]

OpenStudy (anonymous):

Do I have to get common denominators before I add or is the answer 2pi/35?

OpenStudy (anonymous):

@satellite73?

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

the answer is cos(12pi/35)

OpenStudy (anonymous):

use this formula cos(A+B) = cosAcosB - sinAsinB

OpenStudy (anonymous):

@niksva do you mean cos(2pi/35)? if it is cos(12pi/35) how did you get that answer?

OpenStudy (anonymous):

cos(pi/7 + pi/5) = cos(5pi+7pi/35 ) = cos(12pi/35)

OpenStudy (anonymous):

@niksva Thank You!!!!!!